OpenStudy (anonymous):

y varies directly as the square of x. When x = 4, y = 32. Find x when y = 18. x = ±3 x = ±8 x = ±10 x = ±26

OpenStudy (anonymous):

its like this.. \[y \alpha x^2\] => y=kx^2 if x=4 then y=32 put it in above relation to find k i.e. 32=k(4)^2 =>32=16k =>k=2 so if y=18 then.. 18=2x^2 =>x^2=9=>x=+-3 so opton A is right ans.